# Entropic vector quantile regression: Duality and Gaussian case

### v1

@@ 4. Proofs for Section 2
$\begin{split}\varepsilon\mathsf{KL}\left(\pi^{\varepsilon}\,\|\,R\right)&=\varepsilon\int\log\left(\frac{d\pi^{\varepsilon}}{dR}\right)\,d\pi^{\varepsilon}\\ &=\int\big(f^{\varepsilon}(u)+\langle g^{\varepsilon}(u),x\rangle+h^{\varepsilon}(x,y)-c(u,y)\big)\,d\pi^{\varepsilon}(u,x,y)\\ &=\int f^{\varepsilon}\,d\mu+\int h^{\varepsilon}\,d\nu-\int c\,d\pi^{\varepsilon}.\end{split}$
@@
Conclude that
$\begin{split}\mathsf{T}^{\varepsilon}(\mu,\nu)&=\int c\,d\pi^{\varepsilon}+\varepsilon\mathsf{KL}\left(\pi^{\varepsilon}\,\|\,R\right)\\ &=\int f^{\varepsilon}\,d\mu+\int h^{\varepsilon}\,d\nu\\ &=D^{\varepsilon}(f^{\varepsilon},g^{\varepsilon},h^{\varepsilon})\leq\mathsf{D}^{\varepsilon}(\mu,\nu).\end{split}$
@@
By strict convexity of the exponential function, we have
@@
Under Assumption 2.1 (i) (cf. Lemma 4.3), one has $\mathfrak{f}-a=0$ and $\mathfrak{g}-v=0$ $\mu$-a.s. ∎
@@
$\begin{split}0\leq h^{\varepsilon}(x,y_{1})-h^{\varepsilon}(x,y_{2})&=\varepsilon\log\int e^{H(u,x,y_{2})/\varepsilon}\,d\mu(u)-\varepsilon\log\int e^{H(u,x,y_{1})/\varepsilon}\,d\mu(u)\\ &=\varepsilon\log\int e^{H(u,x,y_{1})/\varepsilon}\cdot e^{(c(u,y_{1})-c(u,y_{2}))/\varepsilon}\,d\mu(u)\\ &\qquad-\varepsilon\log\int e^{H(u,x,y_{2})/\varepsilon}\,d\mu(u)\\ &\leq\sup_{u\in\mathcal{U}}(c(u,y_{1})-c(u,y_{2})),\end{split}$
@@
$d\pi_{u_{0}}(x,y)=w_{u_{0}}(x,y)e^{f(u_{0})+\langle g(u_{0}),x\rangle}d\nu_{x}(y)d\kappa(x).$
@@
$\tilde{\kappa}_{u_{0}}(\mathbb{R}^{d_{x}})\leq e^{-f(u_{0})+\|g(u_{0})\|\sup_{x\in\mathcal{X}}\|x\|}<\infty.$
@@
so that $A(\theta,u)=-\|u\|^{2}/2+\log\mathsf{Z}(\theta,u)$. Observe that $\log\mathsf{Z}(\theta,u)$ is the log-partition function corresponding to the exponential family $e^{\langle\theta,x\rangle+\langle u,y\rangle-\log\mathsf{Z}(\theta,u)}$ with base measure $e^{h(x,y)-\|y\|^{2}/2}d\nu(x,y)$ The base measure is finite by a similar argument to the previous proof, and $\mathsf{Z}(\theta,u)$ is finite for all $(\theta,u)\in\mathbb{R}^{d_{x}+d_{y}}$ because $\nu$ is compactly supported. Now, by [BAR14, Theorem 7.2], $\log Z(\theta,u)$ is real analytic on $\mathbb{R}^{d_{x}+d_{y}}$, and so is $A(\theta,u)$.
@@ 5. Proofs for Section 3
Any coupling $\pi\in\Pi(\mu,\nu)$ with $\mathsf{KL}\left(\pi\,\|\,\mu\otimes\nu\right)<\infty$ is absolutely continuous with respect to Lebesgue measure, and the KL divergence decomposes as
$\mathsf{KL}\left(\pi\,\|\,\mu\otimes\nu\right)=\operatorname{Ent}(\mu)+\operatorname{Ent}(\nu)-\operatorname{Ent}(\pi).$
See, e.g., [CNR25, Exercise 4.6.2]. Since $\operatorname{Ent}(\mu)$ and $\operatorname{Ent}(\nu)$ depend only on the marginals, the entropic VQR problem (6) is equivalent to
@@
Next, since $\pi^{\varepsilon}$ and $\pi_{G}$ share the same mean vector and covariance matrix, we have $\mathbb{E}_{\pi_{G}}[\|U-\tilde{Y}\|^{2}]=\mathbb{E}_{\pi^{\varepsilon}}[\|U-\tilde{Y}\|^{2}]$. Furthermore, among all distributions with a fixed mean vector and a fixed covariance matrix, the multivariate Gaussian distribution maximizes the differential entropy. Hence, $\text{Ent}(\pi_{G})\geq\text{Ent}(\pi^{\varepsilon})$. Since $\pi^{\varepsilon}$ is the unique minimizer of (27), we conclude $\pi^{\varepsilon}=\pi_{G}$. ∎
@@
$\begin{split}d\pi^{\varepsilon}(u,x,y)&=\exp\left(\frac{1}{\varepsilon}\left(f^{\varepsilon}(u)+\langle g^{\varepsilon}(u),x\rangle+h^{\varepsilon}(x,y)-\frac{1}{2}\|u-y\|^{2}\right)\right)\,d\mu(u)\,d\nu(x,y)\\ &=\frac{1}{(2\pi)^{(2d_{x}+d_{y})/2}\sqrt{\det(\Sigma)}}\exp\left(\frac{1}{\varepsilon}\left(f^{\varepsilon}(u)+\langle g^{\varepsilon}(u),x\rangle+h^{\varepsilon}(x,y)-\frac{1}{2}\|u-y\|^{2}\right)\right)\\ &\quad\times\exp\left(-\frac{1}{2}\|u\|^{2}-\frac{1}{2}(z-m)^{\top}\Sigma^{-1}(z-m)\right)\,dudz,\end{split}$
@@
Assume without loss of generality that $m_{Y}=0$. We write $w=(u,x,y)\in\mathbb{R}^{N}$ with $N:=d_{x}+2d_{y}$ instead of $w=(u^{\top},x^{\top},y^{\top})^{\top}$ for notational convenience. Observe that $\pi^{o}$ is supported on the linear subspace
@@
Using spectral expansions, one can see that $\Lambda_{\varepsilon}-\Lambda_{0}=O(\varepsilon)$, so that the last two terms on the right-hand side are $O(\varepsilon^{2})$. This finishes the proof of Step 1.

### v2

@@ 4.4.1. Preliminary lemmas
$\begin{split}\varepsilon\mathsf{KL}(\pi^{\varepsilon}\,\|\,R)&=\varepsilon\int\log\left(\frac{d\pi^{\varepsilon}}{dR}\right)\,d\pi^{\varepsilon}\\ &=\int\big(f^{\varepsilon}(u)+\langle g^{\varepsilon}(u),x\rangle+h^{\varepsilon}(x,y)-c(u,y)\big)\,d\pi^{\varepsilon}(u,x,y)\\ &=\int f^{\varepsilon}\,d\mu+\int h^{\varepsilon}\,d\nu-\int c\,d\pi^{\varepsilon}.\end{split}$
@@
Observe that $\iota^{\varepsilon}(f^{\varepsilon},g^{\varepsilon},h^{\varepsilon})=0$ as $\pi^{\varepsilon}$ is a probability measure. Conclude that
$\begin{split}\mathsf{T}^{\varepsilon}(\mu,\nu)&=\int c\,d\pi^{\varepsilon}+\varepsilon\mathsf{KL}(\pi^{\varepsilon}\,\|\,R)\\ &=\int f^{\varepsilon}\,d\mu+\int h^{\varepsilon}\,d\nu\\ &=D^{\varepsilon}(f^{\varepsilon},g^{\varepsilon},h^{\varepsilon})\leq\mathsf{D}^{\varepsilon}(\mu,\nu).\end{split}$
@@
Observe that the dual objective can be written as
$\begin{split}D^{\varepsilon}(f,g,h)&=\int\big(f(u)+\langle g(u),x\rangle+h(x,y)\big)\,dR(u,x,y)\\ &\quad-\varepsilon\int e^{\frac{1}{\varepsilon}(f(u)+\langle g(u),x\rangle+h(x,y)-c(u,y))}\,dR(u,x,y)+\varepsilon\end{split}$
[…]
@@
Under Assumption 2.1 (i) (cf. Lemma 4.5), one has $\mathfrak{f}-a=0$ and $\mathfrak{g}-v=0$ $\mu$-a.e. ∎
@@
$\begin{split}0\leq h^{\varepsilon}(x,y_{1})-h^{\varepsilon}(x,y_{2})&=\varepsilon\log\int e^{H(u,x,y_{2})/\varepsilon}\,d\mu(u)-\varepsilon\log\int e^{H(u,x,y_{1})/\varepsilon}\,d\mu(u)\\ &=\varepsilon\log\int e^{H(u,x,y_{1})/\varepsilon}\cdot e^{(c(u,y_{1})-c(u,y_{2}))/\varepsilon}\,d\mu(u)\\ &\qquad-\varepsilon\log\int e^{H(u,x,y_{1})/\varepsilon}\,d\mu(u)\\ &\leq\sup_{u\in\mathcal{U}}(c(u,y_{1})-c(u,y_{2})),\end{split}$
@@
$d\pi_{u_{0}}(x,y)=w_{u_{0}}(x,y)e^{f(u_{0})+\langle g(u_{0}),x\rangle}d\nu_{x}(y)d\kappa(x),$
that is,
$w_{u_{0}}(x,y)d\nu_{x}(y)d\kappa(x)=e^{-f(u_{0})-\langle g(u_{0}),x\rangle}d\pi_{u_{0}}(x,y).$
@@
$\begin{split}\tilde{\kappa}_{u_{0}}(\mathbb{R}^{d_{x}})&=\int_{\mathbb{R}^{d_{x}+d_{y}}}e^{-f(u_{0})-\langle g(u_{0}),x\rangle}d\pi_{u_{0}}(x,y)\\ &\leq e^{-f(u_{0})+\|g(u_{0})\|\sup_{x\in\mathcal{X}}\|x\|}<\infty.\end{split}$
@@
so that $A(\theta,u)=-\|u\|^{2}/2+\log\mathsf{Z}(\theta,u)$. Observe that $\log\mathsf{Z}(\theta,u)$ is the log-partition function corresponding to the exponential family $e^{\langle\theta,x\rangle+\langle u,y\rangle-\log\mathsf{Z}(\theta,u)}$ with base measure $e^{h(x,y)-\|y\|^{2}/2}d\nu(x,y)$. The base measure is finite by a similar argument to the previous proof, and $\mathsf{Z}(\theta,u)$ is finite for all $(\theta,u)\in\mathbb{R}^{d_{x}+d_{y}}$ because $\nu$ is compactly supported. Now, by [BAR14, Theorem 7.2], $\log Z(\theta,u)$ is real analytic on $\mathbb{R}^{d_{x}+d_{y}}$, and so is $A(\theta,u)$.
@@ 5. Proofs for Section 3
Any coupling $\pi\in\Pi(\mu,\nu)$ with $\mathsf{KL}(\pi\,\|\,\mu\otimes\nu)<\infty$ is absolutely continuous with respect to Lebesgue measure, and the KL divergence decomposes as
$\mathsf{KL}(\pi\,\|\,\mu\otimes\nu)=\operatorname{Ent}(\mu)+\operatorname{Ent}(\nu)-\operatorname{Ent}(\pi).$
Since $\operatorname{Ent}(\mu)$ and $\operatorname{Ent}(\nu)$ depend only on the marginals, the entropic VQR problem (7) is equivalent to
@@
Next, since $\pi^{\varepsilon}$ and $\pi_{G}$ share the same mean vector and covariance matrix, we have $\mathbb{E}_{\pi_{G}}[\|U-\tilde{Y}\|^{2}]=\mathbb{E}_{\pi^{\varepsilon}}[\|U-\tilde{Y}\|^{2}]$. In addition, identifying a measure and its Lebesgue density, we have
$\begin{split}0\leq\mathsf{KL}(\pi^{\varepsilon}\,\|\,\pi_{G})&=\int\pi^{\varepsilon}(w)\log\frac{\pi^{\varepsilon}(w)}{\pi_{G}(w)}\,dw\\ &=-\operatorname{Ent}(\pi^{\varepsilon})+\mathbb{E}_{\pi^{\varepsilon}}\left[-\log\pi_{G}(U,\tilde{X},\tilde{Y})\right]\\ &=-\operatorname{Ent}(\pi^{\varepsilon})+\underbrace{\mathbb{E}_{\pi_{G}}\left[-\log\pi_{G}(U,\tilde{X},\tilde{Y})\right]}_{=\operatorname{Ent}(\pi_{G})},\end{split}$
that is, $\text{Ent}(\pi_{G})\geq\text{Ent}(\pi^{\varepsilon})$. Since $\pi^{\varepsilon}$ is the unique minimizer of (33), we conclude $\pi^{\varepsilon}=\pi_{G}$. ∎
Proof of Theorem 3.1 (i).
@@
$\begin{split}d\pi^{\varepsilon}(u,x,y)&=\exp\left(\frac{1}{\varepsilon}\left(f^{\varepsilon}(u)+\langle g^{\varepsilon}(u),x\rangle+h^{\varepsilon}(x,y)-\frac{1}{2}\|u-y\|^{2}\right)\right)\,d\mu(u)\,d\nu(x,y)\\ &=\frac{1}{(2\pi)^{(d_{x}+2d_{y})/2}\sqrt{\det(\Sigma)}}\exp\left(\frac{1}{\varepsilon}\left(f^{\varepsilon}(u)+\langle g^{\varepsilon}(u),x\rangle+h^{\varepsilon}(x,y)-\frac{1}{2}\|u-y\|^{2}\right)\right)\\ &\quad\times\exp\left(-\frac{1}{2}\|u\|^{2}-\frac{1}{2}(z-m)^{\top}\Sigma^{-1}(z-m)\right)\,dudz,\end{split}$
@@
Assume without loss of generality that $m_{Y}=0$. We write $w=(u,x,y)\in\mathbb{R}^{N}$ with $N:=d_{x}+2d_{y}$ instead of $w=(u^{\top},x^{\top},y^{\top})^{\top}$ for notational convenience. Observe that $\pi^{o}$ is supported on the vector subspace
@@
Using spectral expansions, one can see that $\Lambda_{\varepsilon}-\Lambda_{o}=O(\varepsilon)$, so that the last two terms on the right-hand side are $O(\varepsilon^{2})$. This finishes the proof of Step 1.
@@
Appendix A Partial converse of Lemma 4.1
The following lemma provides a partial converse of Lemma 4.1.
Lemma A.1.
Let $(M,d)$ be a Polish metric space. Suppose that $Q\in\mathcal{P}_{1}(M)$, but
$\int e^{\alpha d(x,x_{0})}\,dQ(x)=\infty$
for every $\alpha>0$ for some $x_{0}\in M$. Then there exists a sequence $P_{n}\in\mathcal{P}_{1}(M)$ such that $\limsup_{n\to\infty}\mathsf{KL}(P_{n}\,\|\,Q)<\infty$ but $\lim_{n\to\infty}\mathsf{W}_{1}(P_{n},Q)=\infty$. In particular, $\mathsf{KL}(\cdot\,\|\,Q)$ is not coercive in $\mathsf{W}_{1}$.
Proof.
Set $\bar{F}(r):=Q(\{d(\cdot,x_{0})\geq r\})$. We first verify that
$\liminf_{r\to\infty}r^{-1}\log\big(1/\bar{F}(r)\big)=0.$
Suppose on the contrary that the left-hand side is strictly positive (including $\infty$). Then, there exist $\beta>0,r_{0}>0$ such that
$\bar{F}(r)\leq e^{-\beta r},\ r\geq r_{0},$
which, however, implies that $d(\cdot,x_{0})$ is sub-exponential under $Q$, contradicting the assumption.
Now, choose $r_{n}\to\infty$ such that
$\lim_{n\to\infty}r_{n}^{-1}q_{n}=0\quad\text{with}\quad q_{n}:=\log\big(1/\bar{F}(r_{n})\big).$
Assuming that $n$ is large enough, we define
$P_{n}:=(1-q_{n}^{-1})Q+q_{n}^{-1}Q_{n}\quad\text{with}\quad dQ_{n}:=e^{q_{n}}\mathbbm{1}_{\{d(\cdot,x_{0})\geq r_{n}\}}\,dQ.$
By convexity of $\mathsf{KL}(\cdot\,\|\,Q)$,
$\mathsf{KL}(P_{n}\,\|\,Q)\leq q_{n}^{-1}\mathsf{KL}(Q_{n}\,\|\,Q)=1.$
On the other hand, by the Kantorovich-Rubinstein duality,
$\begin{split}\mathsf{W}_{1}(P_{n},Q)&\geq\int d(\cdot,x_{0})\,d(P_{n}-Q)=q_{n}^{-1}\int d(\cdot,x_{0})\,d(Q_{n}-Q)\\ &\geq q_{n}^{-1}(r_{n}-a),\end{split}$
where $a:=\int d(\cdot,x_{0})\,dQ<\infty$. Since $\lim_{n\to\infty}q_{n}^{-1}r_{n}=\infty$ by construction, we obtain the desired claim. ∎